# PV-Tuning: Beyond Straight-Through Estimation for Extreme LLM Compression

### v1

@@ 1 Introduction
Recent years have seen the development of ever more capable large language models, attracting immense interest from both researchers and industry. One of the driving factors behind progress in this area is the availability of powerful open LLMs such as Llama [58], Mistral [29, 30], or Phi [36]. The main advantage of open LLMs is that they can be run and fine-tuned locally by end users; however, as state-of-the-art LLMs grow larger, they also become harder to run on commodity hardware. For instance, in order to fit the best available Llama-3 model on a consumer GPU, the model would have to be compressed to below 2 bits per parameter [^1] At the time of writing, the best open model (Llama-3 70B) occupies 130GB in FP16, while most consumer GPUs have 8-24GiB VRAM, some of which must be reserved for the attention cache..
@@ 2 Background
As mentioned above, the two SOTA quantization techniques apply fine-tuning, but only update continuous parameters, such as quantization scales. When optimizing over discrete parameter sets, a standard choice in deep learning is the Straight-Through Estimator (STE) [6, 14]. Prior work on LLM compression proposed to update both continuous and discrete parameters, via STE, both for post-training quantization [65, 55] and for training quantized networks from scratch [27]. However, it was observed early on that STE leads to instability when fine-tuning heavily quantized LLMs [65]. While early results suggest that STE can perform well when training quantized models from scratch [39], this behavior is yet to be validated for highly-performant multi-billion-parameter models, which are the focus of our work.
@@ 4.3 Large-scale Evaluation & Discussion
Fine-tuning efficiency. One limitation of our algorithm is that it requires more compute and memory during the fine-tuning procedure. While 7B models can be fine-tuned on a single GPU, our 70B runs require a server with $1008{\times}A$ 100 or rely on RAM offloading. PV-Tuning shares this drawback with prior methods based on STE [20, 60], as both methods need gradients w.r.t. dequantized weights.
@@ 5 Conclusions
Limitations. We focus our effort on evaluating PV-Tuning with multiple setups and models, but we spent relatively little effort tuning our algorithm for each specific setup. For instance, we always use constant learning rate and $\tau$ with no schedule, and always train on the same data. While this shows robustness of PV tuning, it also means that our results may be improved with better hyperparameters.
@@ Appendix E Straight-through Gradient Estimation
As we discuss in Section 2, straight-through estimation introduces an auxiliary non-quantized weight tensor that is updated using the gradients $\nabla\phi(y)$ w.r.t. quantized weights. The quantized weights are then updated to best approximate this auxiliary buffer, usually in terms of $L2$ error. As a result, if an update to $y-\frac{1}{L}\nabla\phi(y)$ is not large enough to change the parameter, it is still accumulated in a straight-through “buffer”. Eventually, the cumulative effect of several such updates will be large enough that $y^{k}$ will no longer be the solution to Equation (6).
This strategy prevents the algorithm from stalling, but it does so at the cost of convergence guarantees. When applied to extreme LLM quantization (Section 4.2, straight-through estimation initially improves $y^{k}$, but then stops improving and oscillates. To summarize, both stochastic rounding and straight-through estimation can be viewed as ways to address the problem. However, they do so at the cost of changing the optimization landscape and may not converge at all.
@@ Appendix H Calibration Data Matters
To remove the bias, we use standard666from e.g. https://github.com/huggingface/transformers/blob/main/examples/pytorch/language-modeling/run_clm.py LM preprocessing that concatenates all documents, then splits them into evenly sized chunks that become training sequences. The benefits from debiasing range from insignificant to as large as 0.15 perplexity for some models. To compensate for that, we run experiments with the same preprocessing protocol unless explicitly stated otherwise.
@@ Appendix I Additional Engineering Considerations
Both our training runs and almost all baselines use the same sample of RedPajama data from previous sections [^11] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T-Sample. The only exception to this is OneBit. To better explore the effect of training data, we report an additional 13B run at 0.97 bits per parameter using Cosmopedia [5] sample [^12] https://huggingface.co/datasets/HuggingFaceTB/cosmopedia_6M as calibraton data. We tag this run in Table 7 denote that fact that it uses a different finetuning dataset.
@@
* - This (and only this) run uses Cosmopedia [5] dataset instead of RedPajama. This dataset is more similar to the calibraton data described in OneBit.
@@
Table 9: Evaluation of quantized Llama 3 models for 1.x bits per weight, grouped by bitwidth. We report perplexity on WikiText2 [40] & C4 [48] and zero-shot accuracy. The Average is the mean accuracy of 5 zero-shot tasks. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average (Accuracy).
@@
PV-Tuning 2.00 4.58 6.57 50.17 79.00 63.48 79.54 78.06 70.05
Table 10: Evaluation of quantized Mistral v0.1 7B (A) and Phi 3-Mini-4k-Instruct 3.8B (B) models for 1.x and 2.x bits per weight, grouped by bitwidth. We report perplexity on WikiText2 [40] & C4 [48] and zero-shot accuracy. The Average is the mean accuracy of 5 zero-shot tasks. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average (Accuracy).

### v2

@@ 1 Introduction
Recent years have seen the development of ever more capable large language models, attracting immense interest from both researchers and industry. One of the driving factors behind progress in this area is the availability of powerful open LLMs such as Llama [63], Mistral [31, 32], or Phi [38]. The main advantage of open LLMs is that they can be run and fine-tuned locally by end users; however, as state-of-the-art LLMs grow larger, they also become harder to run on commodity hardware. For instance, in order to fit the best available Llama-3 model on a consumer GPU, the model would have to be compressed to below 2 bits per parameter [^1] At the time of writing, the best open model (Llama-3 70B) takes up 130GB in FP16, while most consumer GPUs have 8-24GiB DRAM, some of which must be reserved for the attention cache..
@@ 2 Background
As mentioned above, the two SOTA quantization techniques apply fine-tuning, but only update continuous parameters, such as quantization scales. When optimizing over discrete parameter sets, a standard choice in deep learning is the Straight-Through Estimator (STE) [5, 13, 67]. Prior work on LLM compression proposed to update both continuous and discrete parameters, via STE, both for post-training quantization [72, 58] and for training quantized networks from scratch [29]. However, it was observed early on that STE leads to instability when fine-tuning heavily quantized LLMs [72]. While early results suggest that STE can perform well when training quantized models from scratch [41], this behavior is yet to be validated for highly-performant multi-billion-parameter models, which are the focus of our work.
@@ 4.3 Large-scale Evaluation & Discussion
Fine-tuning efficiency. One limitation of our algorithm is that it requires more compute and memory during the fine-tuning procedure. The 7B models can be fine-tuned on a single GPU, our 70B runs require a server with $1008{\times}A$ 100 or rely on RAM offloading. PV-Tuning shares this drawback with prior methods based on STE [19, 65], as both methods need gradients w.r.t. dequantized weights. Our longest training run took 2 days on 8 GPUs to outperform all baselines and 8 days to fully converge.
@@ 5 Conclusions
Limitations. We focus our effort on evaluating PV-Tuning with multiple setups and models, but we spent relatively little effort tuning our algorithm for each specific setup. For instance, we always use constant learning rate and $\tau$ with no schedule, and always train on the same data. While this shows robustness of PV-Tuning, it also means that our results may be improved with better hyperparameters. Furthermore, the algorithm could achieve better accuracy by simply training longer and on more data.
@@
6 Acknowledgements
Authors would like to thank Vage Egiazarian, Andrei Panferov and Ruslan Svirschevski for their help and advice on AQLM codebase and running large-scale experiments. We also thank Philip Zmushko and Artem Fedorov for helpful discussions during the early stages of our research. Finally, we thank the open-source contributors from llamacpp [^6] https://github.com/ggerganov/llama.cpp and the LocalLlama [^7] https://reddit.com/r/LocalLaMA community for discussions and inspirations on practical use cases of quantized language models.
@@ Appendix E Straight-through Gradient Estimation
Straight-through gradient estimation is a technique for training neural networks with discrete components that ignores these discrete components during backpropagation. Its usage goes back to the Perceptron introduced by Rosenblatt [54]. There, the artificial neuron uses a step function as activation, but the training procedure treats this function as though it was identity. Subsequent works introduce variations of this idea, extend it to multi-layer networks [28], discuss its convergence properties [70, 21, 74]. Other works use straight-through estimation or similar techniques to training neural network with quantized weights [27, 59, 60].
STE for LLM quantization. As we discuss in Section 2, straight-through estimation introduces an auxiliary non-quantized weight tensor that is updated using the gradients $\nabla\phi(y)$ w.r.t. quantized weights. The quantized weights are then updated to best approximate this auxiliary buffer, usually in terms of $L2$ error. As a result, if an update to $y-\frac{1}{L}\nabla\phi(y)$ is not large enough to change the parameter, it is still accumulated in a straight-through “buffer”. Eventually, the cumulative effect of several such updates will be large enough that $y^{k}$ will no longer be the solution to Equation (6).
This strategy prevents the algorithm from stalling, but it does so at the cost of convergence guarantees [74]. When applied to extreme LLM quantization (Section 4.2, straight-through estimation initially improves $y^{k}$, but then stops improving and oscillates. We also tried several a variant of straight-through estimation [60] that introduce stochasticity to forward pass. When applied to extreme LLM quantization, this variant did not diverge like naive STE, but trained much slower and did not reach the same optimum as “deterministic” STE. We attribute this to the fact that adding noise during training can slow down convergence, which also applies to stochastic rounding (Appendix  F).
@@ Appendix H Calibration Data Matters
To remove the bias, we use standard888from e.g. https://github.com/huggingface/transformers/blob/main/examples/pytorch/language-modeling/run_clm.py LM preprocessing that concatenates all documents, then splits them into evenly sized chunks that become training sequences. The benefits from debiasing range from insignificant to as large as 0.15 perplexity for some models. To compensate for that, we run experiments with the same preprocessing protocol unless explicitly stated otherwise.
@@ Appendix I Additional Engineering Considerations
Both our training runs and almost all baselines use the same sample of RedPajama data from previous sections [^13] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T-Sample. The only exception to this is OneBit that uses a corpora of LLM outputs gathered specifically for that paper [71].
@@
Table 9: Evaluation of quantized Llama 3 models for 1-2.3 bits per weight, grouped by bitwidth. We report perplexity on WikiText2 [42] & C4 [50] and zero-shot accuracy. The Average is the mean accuracy of 5 zero-shot tasks. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average (Accuracy).
@@
PV-Tuning 2.07 4.55 6.54 50.77 80.22 63.85 79.22 78.06 70.42
Table 10: Evaluation of quantized Mistral v0.1 7B (A) and Phi 3-Mini-4k-Instruct 3.8B (B) models for 1-2.3 bits per weight, grouped by bitwidth. We report perplexity on WikiText2 [42] & C4 [50] and zero-shot accuracy. The Average is the mean accuracy of 5 zero-shot tasks. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average (Accuracy).